OpenStudy (anonymous):

Type the correct answer in the box. Round off your answer to the nearest integer.

OpenStudy (anonymous):

<A and ∠C are right angles. m∠p = ? degrees.

OpenStudy (anonymous):

@butterflydreamer @cupcakerain @dan815 @Data_LG2 @Demonx341 @dmndlife24 @EclipsedStar @FEARLESS_JOCEY @HelpBlahBlahBlah @IrishBoy123 @january123 @jordanjamesbay @kidrah69 @leonardo0430 @Loser66 @Michele_Laino

OpenStudy (anonymous):

First, find length DB first using pythagorean theorem. Then, apply trig ratio to calculate for angle p :)

OpenStudy (anonymous):

how do you do that @Data_LG2

OpenStudy (anonymous):

\(\sf a^2 + b^2 = c^2\) this is the pythagorean theorem. You are looking for lenght DB, so it will be like: \(\sf DA^2 + AB^2 = DB^2\) plug in your given values and solve for DB

OpenStudy (anonymous):

so 61?

OpenStudy (anonymous):

don't forget to take the square root of that to remove the exponent of DB

OpenStudy (anonymous):

so 3721

OpenStudy (anonymous):

no, "square root" \(\sf \sqrt{61}\) <---- like this now, do you know trig ratios? SOH CAH TOA?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

 you are given the adjacent and the hypotenuse. Which trig ratio are you going to use to solve for angle p?

OpenStudy (anonymous):